Suppose a random variable X has support {1,2,3,4,5} and that P(X = 1) = P(X = 2) =
2P(X = 3) = 2P(X = 4) = 3P(x = 5). Find the distribution of X and E[X]. .
Answer #1

P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)=1

P(X=1)+P(X=1)+P(X=1)/2+P(X=1)/2+P(X=1)/3=1

(6P(X=1)+6P(X=1)+3P(X=1)+3P(X=1)+2P(X=1))/6=1

20P(X=1)/6=1

P(X=1)=6/20=0.3

Distribution of X is

X. P

1. 0.3

2. 0.3

3. 0.15

4. 0.15

5. 0.1

E(X)